3.6 grammes of magnesium combined with 1.4 grammes of nitrogen. Calculate the simplest formula of the compound formed
1
Expert's answer
2019-10-02T08:02:02-0400

Moles of "Mg=\frac{3.6}{24}=0.15mol"

Moles of "N_2=\frac{1.4}{28}=0.05mol"

As number of moles of "Mg" reacted"=3\times" number of moles of "N_2" reacted

Hence the simplest formula of the compound formed is "Mg_3N_2"